A consumer has a utility function given by lnU =5linx subscrip1 +3lnx subscript 2 if the budget constraint is given by 10x subscript 1 +14x subscript 2 =124 find the optimal quantity of the two commodities
1
Expert's answer
2021-07-26T23:00:01-0400

Need to find-

optimal quantity of the two commodities

Given in the question-

"lnU =5lnx_1 +3lnx_2"


"lnU = lnx_1^5 + lnx_2^3"


"lnU = ln(x_1)^5\times(x_2)^3"


"U = (x_1)^5\times(x_2)^3.................(1)"


Budget equation-

"10x_1 + 14x_2 = 124 ...............(2)"


Equilibrium condition -

"\frac{MU_x}{MU_y} =\frac{ P_1}{P_2}"


"MU_x= 5\times(x_1)^4\times(x_2)^3"


"MU_y = 3(x_1)^5\times(x_2)^2"


"\frac{MU_x}{MU_y }= \frac{(5\times(x_1)^4\times(x_2)^3)}{(3(x_1)^5\times(x2)^2)}"


"\frac{MU_x}{MU_y} =\frac{ 5x_2}{3x_1}"


"\frac{P_1}{p_2}= \frac{10}{14} = \frac{5}{7}"


"\frac{5x_2}{3x_1} =\frac{ 5}{7}"


"7x_2 = 3x_1"

Using this relation and putting the in equation "2"

So,

"x_1 = 7"

"x_2 = 3"

Answer -

Optimal bundle will be (7,3)